Top at the SSC and LHC

P. Sinervo University of Toronto November 6, 1992

I Introduction
II Top Production and Decay
III Detector Requirements
IV Mass Measurement
V Branching Ratios
VI Charged Higgs Complications
VII Associated Production of Top
VIII Conclusions

Introduction

The standard model requires a top quark:

• asymmetry in e+ e− → b b scattering

• limits on FCNC in b decays

• B∘ − B̄∘ mixing

• Electroweak radiative corrections Mtop ∼ 130 ± 35 GeV/c2

Studying Top at the SSC/LHC

Strong motivations to study top at SSC and LHC

• CDF/D0 will only have small top samples, e.g. – ∼ 100 e±μ∓ events for 500 pb−1and Mtop= 130 GeV/c2 – Measure mass to ∼ 5 GeV/c2

• Intrinsically large σt t in pp Collisions – Measure mass and branching ratios

– Measure couplings and other properties

– Search for new physics

• Significant background to Higgs search – H → W + W − vs t t → W + b W −b

– Intrinsic S/N is O(10−3)

• Top “tag” may be useful/essential in Higgs search

pp → H t t

|_{⟶}γγ

Top Production and Decay

σ_{t t} has been calculated to order α_{s}^{3} by Nason et al.

(NP B303, 607 (1988)).

• Uncertainties are order ±30%. (NP B308, 724 (1988))

Rates are enormous by Tevatron standards

SSC 12 nb 1.2 × 107 per year LHC  2 nb 2.0 × 107 per year  for Mtop= 150 GeV/c2

Various processes contribute to top production

g g → t t
q q → t t
W g → t b̄
g b → t W

Associated production is also important for Higgs search

pp → H t t

The standard model top decays via charged current:

• semi-leptonically:

t → W + b

|_{⟶} \{ \begin{matrix}e+ νe \\ μ+ νμ \\ τ+ ντ \end{matrix} BR = 1/3

• hadronically:

t → W + b

|_{⟶} \{ \begin{matrix}c s̄ \\ u d̄ \end{matrix} BR = 2/3

Other modes are predicted in non-minimal models:

• E.g., if a charged Higgs exists, then also could have

t → H+ b
|_{⟶} \{ \begin{matrix}c s̄ \\ τ+ ντ \end{matrix}

Top quark typically produced with < P_{T} >∼ M_{top}/2

• decay products are at moderate PT

• and pseudorapidity |η| < ∼ 2.5

Decay modes involving leptons are most promising

• Reduces QCD backgrounds

Cf. UA1, UA2 and CDF

• Provides a signature for charged current decay

• Leptons are isolated and have PT > ∼ 20 GeV/c

Primary backgrounds come from W + jet production

Detector Requirements

1. Identify isolated, moderate P_{T} leptons

• (> ∼ 30 − 40 GeV/c) gives several Hz of leptons • requires fine segmentation (Δη × Δϕ ∼ 0.05 × 0.05)

2. Reconstruct complex jet topologies

• Calorimeter resolution of σE ∼ 0.7√E

• Hadron calorimeter segmentation of ∼ 0.1

3. Measure E̸T reasonably well

4. Tag b hadrons • Use non-isolated leptons or displaced vertices

• Perhaps essential for lepton + jets channel

• Useful for background studies

Lepton isolation is an important tool to reduce backgrounds

• Energy in cone of radius R ≡ √Δη2 + Δϕ2 = 0.3

MC simulation of physics and electronics pileup show that isolation is not seriously compromised

• Somewhat more sensitive at LHC

• Can use long shaping times for calorimeter readout

Dilepton Channels

The e^{±} μ^{∓} channel is the cleanest channel

• Rate is equal to e+ e− and μ+ μ− combined

• No Drell-Yan backgrounds

At SSC, requiring lepton P_{T}

• PT > 20 GeV/c

• |η| < 2.5

yields ∼ 10^{6} events per SSC year (10^{40} cm^{−2})

Theoretical uncertainty in σ_{t t}(M_{top}) ⇒ σ_{Mtop} ∼ 10 − 15 GeV/c^{2} A similar analysis for LHC illustrates cleanliness

• Require lepton PT > 50 GeV/c

• ∑ ET < 10 GeV in cone of R = 0.4 for both leptons • |η| < 1.5

yields 3 × 10^{5} events in ∫ L dt = 10^{41} cm^{−2} A more sensitive M_{top} measurement comes from sequential dileptons

t → W^{+} b → μ^{−}X
\underrightarrow{|}e^{+} ν_{e}

Identify W → e^{+} ν_{e} and b → μ^{−}X using isolation

• Electron PT > 20 GeV/c

• ∑ ET < 4 Gev in R < 0.2 around electron

• Muon PT > 20 GeV/c

• ∑ ET > 20 Gev in R < 0.4 around muon

• PT (eμ) > 100 GeV/c

⇒ 7 × 10^{4} events per year < M(eμ) > sensitive to M_{top} Uncertainty in M_{top} dominated by systematic uncertainties

Effect  σMtop (GeV/c2)
b quark fragmentation  1.5
t quark PT  1.9
Statistics (1 year)  0.5

⇒ Can expect to measure M_{top} to ∼ 2 − 3 GeV/c^{2}

Lepton + Jets

The lepton + jets mode provides inherently more rate

• Can perform full reconstruction

• Can directly measure Mtop and BR

• However, jet combinatorics are a problem

SSC and LHC groups have made extensive studies

• Agree that kinematical selection can reconstruct top – Require isolated lepton with PT > ∼ 40 GeV/c – Require 3 jets with high PT cut

• More efficiently done using b quark tag

SDC approach takes advantage of b quark tagging

• Allows measurement to be done at low luminosity

• Better control of systematics in Mtop measurement

Selection requires

• Isolated lepton with PT > 40 GeV/c

• ≥ 3 jets (R = 0.5 cone used for clustering)

• PT > 30 and |η| < 2.5 for jets

• One jet to be tagged as b quark (ϵ ∼ 30%)

• PT > 200 GeV/c of 3-jet system

Resulting event rate for 1 SSC year is M(jj) distribution for non-b jets shows clear W

• Require 65 < M(jj) < 95 GeV/c2

Systematics in M_{top} dominated by jet energy scales

• Use MC to correct for clustering effects, etc.

• Can calibrate low-energy jet scale using W signal

• b quark fragmentation

σ_{Mtop} ∼ 2 − 3GeV/c^{2}

Top Branching Ratios

Although e^{±} and μ^{±} decay modes are most accessible,

t → τ ντ b

is detectable using single-prong τ → π^{−}X decays

Require

• isolated lepton with PT > 40 GeV/c(from W decay)

• one jet with PT > 30 tagged as b hadron

• isolated, charged track with PT > 50 GeV/c

⇒ uncertainty in relative τ rate is < ∼ 5% Observation of top in

pp → e± μ∓  ∝ BR(t → e νe b) × BR(t → μ νμ b)
pp → l± l∓  ∝ BR(t → l νl b)2
pp → μ± + jets  ∝ BR(t → μ νμ b) × BR(t → q q̄ b)
pp → e± + jets  ∝ BR(t → e νe b) × BR(t → q q̄ b)
pp → τ± τ∓  ∝ BR(t → τ ντ b) × BR(t → q q̄ b)

Allows measurement of relative BR’s to 1 − 5%

• Limited by systematic uncertainty in efficiency

• Depends on isolation of lepton and physics models

Charged Higgs

Heavy charged Higgs coupling to top is detectable

• May see it in BR(t → τX)/BR(t → e/μX)

Can also reconstruct it’s 3-jet final states in pp → t t → W^{+} b H^{−}b̄

• Follow same jet reconstruction technique

• Cut on three-jet system to reduce combinatorics

• Require one b quark jet be tagged

Resulting M(jj) distribution for M_{H+} = 125 GeV/c^{2} is

22

H^{+} is clearly detectable over a large range of tan β.

• Characterize signal as excess over background

• Plot statistical significance vs tan β

Associated Top Production

pp → H t t may be important for a light Higgs search

• The W W channel is closed and Z Z∗ is rate-limited

• H → γγ is largest potential rate – QCD diphoton backgrounds are enormous

Tag t t system with isolated, high P_{T} lepton

• Reduces QCD fake backgrounds to manageable level

• Isolation removes bb̄ γγ background

• No longer need high-precision EM calorimetry

“Reach” in γγ l X channel complements Z Z^{∗}

25

Background to Higgs Search

Extensive studies have looked at implications of “heavy” top to Higgs search

• Forced to abandon “simple” H → W + W − search

• Requires substantially more sophisticated approach – W W fusion process and forward jet tagging – stringent kinematic cuts

Makes understanding t t production and decay kinematics even more important

Conclusions and Summary

SSC and LHC are inherent “top factories”

• Dominate low and intermediate mass scale physics

• Next generation “calibration” signal

Measure M_{top} to ∼ 2−3 GeV/c^{2} in 1 year of operation

• Limited by systematics

• Could eventually reduce uncertainty to < ∼ 1 GeV/c2(?!)

Branching ratios measured to < ∼ 5% in 1 year

• Test universality of couplings

• Sensitive to charged Higgs

Useful tag for low mass Higgs (M_{H} < ∼ 100 GeV/c^{2})

Serious background to H → W^{+} W^{−}